OpenStudy (anonymous):

find the edge of a cube if its total surface area is numerically equal to its volume. Guys answer it please . :* mwah.

OpenStudy (anonymous):

let x be the edge of a cube. surface area=6*(x)^2=6x^2 volume=x^3 x^3=6x^2 or x^3-6x^2=0 can you solve it ?